OpenStudy (anonymous):

53) The measure of a supplement of an angle is 30 times more than 5 times the measure of an angle. Find the measure of the angle. a) 15 deg. b) 25 deg. c) 65 deg. d) 115 deg.

OpenStudy (anonymous):

30 times more than or 30 degrees more than...?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 5x + 30 + x = 180 }\) \(\Large \color{purple}{\rightarrow 6x = 150 }\) \(\Large \color{purple}{\rightarrow x = 25 }\) So, 25 degrees and 155 degrees are the answers respectively(Sum of supplements = 180)

OpenStudy (anonymous):

yup. 30 times more than 5 times the measure of an angle

OpenStudy (anonymous):

okay, sum of an angle and its supplement=180 degree say, the angle is x so, its supplement=5x+30 5x+30+x=180 x=25 degree (b)

OpenStudy (anonymous):

30 times more..?? doesn't make any sense to me.. :/

OpenStudy (anonymous):

so the answer is 25 deg?

Parth (parthkohli):

It is, yes.

OpenStudy (anonymous):

i guess so

OpenStudy (lgbasallote):

my interpretation: if it's 30 times more that 5 times the angle then.... i see 2 more possible solutions 5x + 30x + x = 180 OR 30(5+x) + x = 180 your choice :D

OpenStudy (lgbasallote):

the second equation is 30(5x) + x = 180